Translate the following mathematical statements into mathematical equations and solve the unknown
1. The amount raised during JS Promenade Night varies directly as the number of attendees and inversely as the amount of expenses used on the said event. If the amount of money raised for 250 attendees was Php5000.00 and the amount for expenses was Php4000.00 How much money will bhe raised for 300 attendees and Php6000.00 amount of expenses used during the JS Promenade Night?
2. In particular school, the number of girls varies directly as the number of boys and inversely as the number of teachers. When there were 40 girls, there were 25 teachers and 10 boys. How many boys when there were 30 girls and 60 teachers?

Answer by Boreal(15235)

M=k*A/E, Money is M, A attendees, E expenses
5000=k*250/4000
k=5000*4000/250=80000
-
M=80000A/E=80000*300/6000
M=4000
Another way of looking at it is the attendees for the second is 1.2 x that of the first and the cost is 1.5 times that of the first.
1.2/1.5=0.8
5000*0.8=4000Php